OpenStudy (anonymous):

-6y=-3x+2 -3y=6x-1 parallel perpendicular neither

OpenStudy (texaschic101):

put your equations in y = mx + b form and compare the slopes(m). -6y = -3x + 2 y = 1/2x - 1/3 slope = 1/2 -3y = 6x - 1 y = -2x + 1/3 slope = -2 parallel lines will have same slope. perpendicular lines will have negative reciprocal slopes....all that means is " flip " the slope and change the sign. 1/2 and -2 are negative reciprocals...these lines are perpendicular

OpenStudy (anonymous):

oh and thank you :)